Find the x-intercepts of the graph of the equation:
y = x^2 - 1
i really need your help asap thanx :)

Answer by ewatrrr(24785)


Hi,
Finding the x-intercepts of the graph of the following equation
y = x^2 - 1 [parabola crosses x-axis when y = 0
0 = x^2 - 1
x^2 = 1
x = ± 1 x-intercepts are points Pt(-1,0) and Pt(1,0)